2. Step 1: A study of iron deficiency among infants compared samples of infants following different feeding regimens. One groStep 2: A study of iron deficiency among infants compared samples of infants following different feeding regimens. One groupStep 3: A study of iron deficiency among infants compared samples of infants following different feeding regimens. One groupStep 4: A study of iron deficiency among infants compared samples of infants following different feeding regimens. One group
Answer #1

Step 1

The given data is summarized as follows:

Sample 1 Sample 2
Sample size n1=23 n2=19
Sample mean 11=13.3 \bar{x}_2=12.4
Sample SD s1=1.7 s2=1.8

The test statistic is given by

t=\frac{\bar{x}_1-\bar{x}_2}{\sqrt{\frac{s_1^2}{n_1}+\frac{s_2^2}{n_2}}}

\Rightarrow t=\frac{13.3-12.4}{\sqrt{\frac{1.7^2}{23}+\frac{1.8^2}{19}}}=\frac{0.9}{0.5442228309}=1.653734369\simeq 1.654

The df is obtained as

df=\frac{(\frac{s_1^2}{n_1}+\frac{s_2^2}{n_2})^2}{\frac{(s_1^2/n_1)^2}{n_1-1}+\frac{(s_2^2/n_2)^2}{n_2-1}}

\Rightarrow df=\frac{(\frac{1.7^2}{23}+\frac{1.8^2}{19})^2}{\frac{(1.7^2/23)^2}{23-1}+\frac{(1.8^2/19)^2}{19-1}}=\frac{0.08772169776}{0.002333170139}=37.59764293\simeq 37.6

The test statistic is 1.654 and the df is 37.6

The p-value is obtained as 0.053227

As the p-value is more than 0.03, we reject the null hypothesis at 3% level of significance.

Step 2

Thus it can be concluded that breast fed children have higher levels of blood haemoglobin at 0.03 significance level. The correct option is the first option

Step 3

The 95% confidence interval is given by

((\bar{x}_1-\bar{x}_2)\pm \sqrt{\frac{s_1^2}{n_1}+\frac{s_2^2}{n_2}}\;t_{0.025,37.6})

i.e.\;((13.3-12.4)\pm \sqrt{\frac{1.7^2}{23}+\frac{1.8^2}{19}}*2.025102)

i.e.\;(0.9\pm 1.102106743)

i.e.\;(-0.202106743,2.002106743)

i.e.\;(-0.202,2.002)

So the correct option is the first option

Step 4

The required assumptions are that the observations must be independent and the parent distribution must be Normal. So the correct option is the first option

Hopefully this will help you. In case of any query, do comment. If you are satisfied with the answer, give it a like. Thanks.